A Complete Bibliography of Metron (2020–2029)

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

       E-mail: beebe@math.utah.edu, beebe@acm.org,
                  beebe@computer.org (Internet)
      WWW URL: http://www.math.utah.edu/~beebe/

17 April 2024 Version 1.08

Title word cross-reference

∘ [51].

19 [29].

absence [13]. absolute [39]. accuracy [63]. activity [58]. Adaptive [93, 69, 30]. administrative [92]. Advances [85, 32]. affect [94, 97]. affluence [11]. against [94]. algorithm [64]. analyses [16]. Analysis [65, 90, 23, 41, 54, 53]. application [42, 93]. applications [82, 85, 91, 49, 47]. applied [66]. Approach [29, 76, 30, 95, 75]. arbitrary [81]. archives [59]. area [96, 99]. assessing [62]. association [52]. asymptotically [26]. attitude [57]. Augmenting [95]. autoregressive [25].

based [11, 78]. Bayes [37]. Bayesian [25, 22, 18, 83, 46, 29]. behaviour [60]. being [58]. Benefit [86]. Benefit-of-the-Doubt [86]. beta [81]. between [83, 89]. biased [83]. binary [71]. binomial [70, 34]. bivariate [81]. BoD [86]. Bonferroni [12]. book [49]. Boundary [43]. breakpoints [46]. breaks [25]. Brq [22]. Building [59]. business [95].

capacity [6]. cardinal [13]. case [94]. censored [64]. census [92, 93]. century [10]. challenge [33]. child [14]. choice [80]. cluster [71]. Clustering [23]. Cochran [24]. collection [57]. combination [63]. combining [95]. Commentary [3, 4, 7, 48, 8]. comparability [13]. comparative [90, 84]. comparison [70, 89]. comparisons [83]. competing [37, 68]. completion [83]. component [61, 69]. component-wise [69]. components [35, 31]. composite [95]. composition [21]. concentration [19]. conditional [17]. confidence [66, 70, 26]. Considerations [49]. considering [20]. constraints [86]. construction [65]. consumer [56]. contingency [42, 52]. continuity [71]. continuous [65]. contributions [33]. copula [65]. Correction [47, 54, 99, 48, 98, 71]. correlation [81]. count [73]. counts [5, 78]. COVID [29]. COVID-19 [29]. credible [70]. credit [80]. cumulative [11, 68]. curve [19].

Data [72, 98, 91, 42, 17, 92, 35, 27, 94, 74, 56, 60, 77, 95]. de-biased [83]. definite [66]. definition [19]. definitions [82]. density [17, 43]. dependence [11]. deprivation [11]. design [92, 93]. designs [71]. detection [36, 46]. diagonal [11]. diagram [11]. Differentially [27]. dimensional [36]. directional [42]. discriminant [16]. dissimilarity [62]. distance [20, 62]. distribution [47, 64, 84, 81]. distributions [66, 21, 69, 6, 79]. does [10]. dominance [21]. Doubt [86]. drawbacks [31]. driven [76]. drivers [89]. dual [76]. dual-frame [76]. Dynamic [88].

editor [47]. Editorial [85, 32]. educational [13]. effect [54, 53]. effects [96, 99]. Efficient [73, 26]. Egypt [14]. election [27]. electric [90]. Eliazar [40]. employment [75]. equality [13]. ergodic [17]. Estimate [44]. estimating [64]. estimation [25, 96, 99, 34, 43, 76, 45, 26, 37, 38, 81, 31]. estimator [43]. estimators [83]. Europe [87]. evaluate [57]. Evaluating [89]. Evaluation [58, 51, 55, 80]. events [41]. example [70]. examples [82]. expectations [88]. exponential [47].

factor [23, 54, 53]. fair [70]. Fay [97]. fixed [20, 26]. fixed-width [26]. Flexible [5]. forecasting [63]. Foreword [72, 98, 91, 9]. frame [76]. framework [78]. Friuli [58]. frontiers [51]. fully [78]. function [19]. functional [17, 35]. functions [68]. fuzzy [43]. FWCI [26].

GARCH [44]. gather [94]. Gaussian [82]. Gender [87, 54, 53]. gender-effect [54, 53]. general [26]. generalised [21]. Generalized [16, 55]. Gini [49, 16, 12, 6]. Giorgi [50]. Giovanni [50]. Giulia [58]. global [42, 52]. group [54, 53, 62]. GWPR [89]. GWR [89].

Haenszel [24]. Handling [61]. Herriot [97]. heterogeneity [41, 89, 62]. Hierarchical [29]. high [5, 36]. high-dimensional [36]. high-inflation [5]. higher [54, 53]. higher-order [54, 53]. housing [92, 46]. hundred [1]. hyperbolic [44].

Iddo [40]. If [94]. Impact [29]. Imposing [86]. Improving [92]. INARCH [34]. incidence [68]. income [20, 6]. increase [60]. index [49, 6]. indicator [95]. indicators [18]. indices [39, 12]. inequalities [14, 87]. Inequality [15, 2, 49, 3, 4, 39, 12, 20, 7, 48, 8]. inequality-indices [12]. Inference [77, 79]. inflation [5]. information [94, 95]. innovation [89]. Integrated [79]. Integration [72, 98, 91, 77]. inter [62]. inter-group [62]. interactions [23]. interval [26]. intervals [70]. Investigating [41]. issue [85, 72, 98, 91, 32]. Italian [92, 93]. Italy [75]. item [80, 73].

Job [54, 53]. judgments [88].

Kruskal [24].

Laws [40]. learning [75]. least [31]. Lehmann [31]. Letter [47]. likelihood [79]. Lindley [47]. Lindley-exponential [47]. linear [23, 17, 16, 97, 46]. lists [73]. literature [90]. Local [36, 17]. lognormal [66]. long [28]. longitudinal [59, 35, 18, 65]. low [83]. low-rank [83]. Lp [63]. Lp-norm [63].

M [44]. M-Estimate [44]. machine [75]. main [33]. malnutrition [14]. Mantel [24]. mapping [77]. marginals [81]. Maria [50]. Marks [29]. massive [92]. matrices [66]. matrix [83]. mean [26, 24]. means [66]. measure [42]. measures [20]. Measuring [13, 11]. media [60]. meta [41]. meta-analysis [41]. methodology [96, 99]. Methods [72, 98, 91, 90, 49, 85]. metric [62]. Metron [1, 10]. Metropolis [69]. Mind [96, 99]. Minimax [45]. minimum [26]. Missing [60, 74]. mixtures [66]. mobility [90]. model [25, 28, 18, 97, 86, 46, 67]. Modeling [74, 29]. modelization [17]. Modelling [56, 85, 88]. models [44, 34, 5, 57, 56, 45, 38, 68]. monograph [40]. most [90]. Movimento [58]. MRPE [26]. Mukhopadhyay [49]. Multi [75, 54, 53]. multi-group [54, 53]. Multi-source [75]. multicollinearity [61]. multidimensional [2, 3, 4, 7, 48, 8]. multilevel [57]. multinomial [79]. multiple [25, 96, 99, 69, 80]. multiple-choice [80]. multiple-try [69]. Multivariate [82, 96, 99, 64, 38].

negative [34]. networked [74]. Nitis [49]. Non [68, 23, 77]. non-linear [23]. Non-parametric [68]. non-probability [77]. norm [63]. normal [64, 26]. notes [40].

Obituary [50]. obtained [21]. odds [52]. official [91]. old [10]. one [73]. opportunity [13]. Optimal [63]. optimality [26]. order [54, 53, 26]. Ordinal [67, 52, 56, 65]. outcomes [71]. outlier [36].

package [22]. pair [65]. panels [59]. paradox [39]. Parameter [81]. parameters [64]. parametric [21, 68]. Pareto [84]. Partha [49]. partial [80]. partition [27]. passenger [55]. people [57]. perceptions [56]. permanent [92]. physical [58]. piecewise [46]. Plateau [69]. point [26]. polytomous [67]. population [92, 78, 93]. populations [74]. positive [66, 28, 81]. Power [40]. Pratim [49]. prediction [30]. predictor [97]. preferences [18]. presence [37]. price [88]. prices [46]. principal [35, 61]. private [27]. probability [77]. processes [82]. program [58]. progress [13]. projections [36]. promoting [58]. properties [47]. proportions [70]. proposal [69]. public [56]. Publisher [99]. purely [26].

quadratic [16]. quality [94, 18, 56]. quantifying [18]. quantile [22, 61].

R [22]. random [28, 96, 99]. randomized [71]. rank [83]. ranking [62]. rare [41]. Rasch [67]. ratios [52, 15]. real [91]. recurrent [68]. regarding [18]. regions [66]. register [78]. regression [22, 28, 61, 5]. regressive [20]. related [90]. relations [21]. relative [39, 20]. release [27]. relevant [19]. represent [6]. representativity [59]. respondent [76]. respondent-driven [76]. response [67]. responses [64, 65]. reverse [42]. review [2, 3, 4, 7, 48, 8]. risk [26]. risks [37, 68]. Robust [35, 37, 32, 38, 34, 33]. Robustness [66]. ROC [19]. RSS [71]. RSS-structured [71].

sample [93]. sampled [74]. samples [76, 77]. sampling [97, 30, 93]. satisfaction [55, 54, 53]. Scheffé [31]. science [33, 77]. score [24]. scoring [80]. search [90]. second [26]. second-order [26]. Sengupta [49]. Sensitivity [20]. separate [57]. sequential [26]. sequentially [27]. series [45]. service [56]. set [43]. show [10]. signal [60]. Simulation [83]. skew [64]. skew-normal [64]. skewed [28]. small [96, 99]. social [60, 93]. Socioeconomic [14]. Some [40, 90]. source [75]. Sources [72, 98, 91]. Southern [87]. sparse [35]. spatial [89]. special [85, 72, 98, 91, 32]. square [42]. squares [31]. stationary [44]. statistic [55]. Statistical [51, 40, 72, 98, 91, 78, 27, 56, 85, 33]. Statistics [32, 91, 33, 95, 75]. status [75]. Stochastic [21]. strategies [26]. structural [25]. structured [71]. Student [29, 18]. studies [41]. study [84, 77]. studying [54, 53]. Survey [72, 98, 91, 94, 77]. surveys [93]. symmetric [66]. symmetry [42]. systems [51, 56].

tables [42, 52]. Taguchi [55]. tail [28]. Teacher [54, 53]. teaching [18]. techniques [51, 73]. technologies [51]. test [80, 68]. Testing [84]. tests [24]. textual [95]. Theoretical [80]. threshold [25]. time [45]. tools [94, 91]. topics [90]. traditional [95]. transfers [20]. transition [92]. transport [56]. trek [12, 40]. try [69]. Tsallis [12]. two [52, 73]. two-way [52]. type [84].

unaided [42]. unidimensional [2, 3, 4, 7, 48, 8]. unsupervised [86]. urban [56]. use [92, 61]. used [90]. Using [57, 63, 56, 65].

value [60]. variables [28]. variance [31]. variances [97]. variation [67]. variational [37]. Venezia [58]. versus [39]. via [27]. violence [94]. vision [42].

Wallis [24]. waste [57]. way [52]. weighted [62]. well [58]. well-being [58]. widely [90]. width [26]. wise [69]. women [94]. work [87]. wrapped [38].

years [1]. young [57].

zeros [5].

References

Alfo:2020:HYM

[1] Marco Alfó. A hundred years of Metron. Metron, 78(1):1–3, April 2020. CODEN MRONAM. ISSN 0026-1424 (print), 2281-695X (elec-

    tronic). URL http://link.springer.com/article/10.1007/s40300-
    020-00166-6;   http://link.springer.com/content/pdf/10.1007/s40300-
    020-00166-6.pdf.

Andreoli:2020:UMI

[2] Francesco Andreoli and Claudio Zoli. From unidimensional to multidimensional inequality: a review. Metron, 78(1):5–42, April 2020. CODEN MRONAM. ISSN 0026-1424 (print), 2281-695X (electronic). URL http:/

    /link.springer.com/article/10.1007/s40300-020-00168-4. See com-

mentary [3, 4, 7, 8].

Arnold:2020:CUM

[3] Barry C. Arnold. Commentary on “From unidimensional to multidimensional inequality: a review”. Metron, 78(1):43–46, April 2020. CODEN MRONAM. ISSN 0026-1424 (print), 2281-695X (electronic). URL http:/

    /link.springer.com/article/10.1007/s40300-020-00169-3. See [2].

Bosmans:2020:CUM

[4] Kristof Bosmans. Commentary on “From unidimensional to multidimensional inequality: a review”. Metron, 78(1):47–50, April 2020. CODEN MRONAM. ISSN 0026-1424 (print), 2281-695X (electronic). URL http://

link.springer.com/article/10.1007/s40300-020-00167-5;   http:
//link.springer.com/content/pdf/10.1007/s40300-020-00167-5.pdf.

See [2].

Goncalves:2020:FRM

[5] Jussiane Nader Gonçalves and Wagner Barreto-Souza. Flexible regression models for counts with high-inflation of zeros. Metron, 78(1):71– 95, April 2020. CODEN MRONAM. ISSN 0026-1424 (print), 2281-

695X (electronic). URL http://link.springer.com/article/10.1007/
s40300-020-00163-9.

Liu:2020:CGI

[6] Yang Liu and Joseph L. Gastwirth. On the capacity of the Gini index to represent income distributions. Metron, 78(1):61–69, April 2020. CODEN MRONAM. ISSN 0026-1424 (print), 2281-695X (electronic). URL http:/

/link.springer.com/article/10.1007/s40300-020-00164-8.

Mosler:2020:CUM

[7] Karl Mosler. Commentary on “From unidimensional to multidimensional inequality: a review”. Metron, 78(1):51–54, April 2020. CODEN MRONAM. ISSN 0026-1424 (print), 2281-695X (electronic). URL http:/

/link.springer.com/article/10.1007/s40300-020-00165-7. See [2].

Weymark:2020:CUM

[8] John A. Weymark. Commentary on “From unidimensional to multidimensional inequality: a review”. Metron, 78(1):55–59, April 2020. CODEN MRONAM. ISSN 0026-1424 (print), 2281-695X (electronic). URL http:/

/link.springer.com/article/10.1007/s40300-020-00170-w. See [2, 7,

48].

Silber:2020:F

[9] Jacques Silber. Foreword. Metron, 78(2):97–98, August 2020. CODEN MRONAM. ISSN 0026-1424 (print), 2281-695X (electronic). URL https:

//link.springer.com/article/10.1007/s40300-020-00181-7.

Giorgi:2020:MCO

[10] Giovanni Maria Giorgi. Metron, a century old, but it does not show it. Metron, 78(2):99–101, August 2020. CODEN MRONAM. ISSN 0026-1424 (print), 2281-695X (electronic). URL https://link.springer.com/

article/10.1007/s40300-020-00182-6.

Decancq:2020:MCD

[11] Koen Decancq. Measuring cumulative deprivation and affluence based on the diagonal dependence diagram. Metron, 78(2):103–117, August 2020. CODEN MRONAM. ISSN 0026-1424 (print), 2281-695X (electronic). URL

https://link.springer.com/article/10.1007/s40300-020-00173-7.

Eliazar:2020:GBT

[12] Iddo Eliazar and Giovanni M. Giorgi. From Gini to Bonferroni to Tsallis: an inequality-indices trek. Metron, 78(2):119–153, August 2020. CODEN MRONAM. ISSN 0026-1424 (print), 2281-695X (electronic). URL https:

//link.springer.com/article/10.1007/s40300-020-00171-9.

Anderson:2020:MPE

[13] Gordon Anderson, Maria Grazia Pittau, and Roberto Zelli. Measuring the progress of equality of educational opportunity in absence of cardinal comparability. Metron, 78(2):155–174, August 2020. CODEN MRONAM. ISSN 0026-1424 (print), 2281-695X (electronic). URL https://link.springer.

com/article/10.1007/s40300-020-00172-8.

Abu-Ismail:2020:SIC

[14] Khalid Abu-Ismail, Verena Gantner, and Myra Yazbeck. Socioeconomic inequalities in child malnutrition in Egypt. Metron, 78(2):175–191, August 2020. CODEN MRONAM. ISSN 0026-1424 (print), 2281-695X (elec-

tronic). URL https://link.springer.com/article/10.1007/s40300-
020-00176-4.

Yalonetzky:2020:IR

[15] Gaston Yalonetzky. Inequality of ratios. Metron, 78(2):193–217, August 2020. CODEN MRONAM. ISSN 0026-1424 (print), 2281-695X (elec-

tronic). URL https://link.springer.com/article/10.1007/s40300-
020-00177-3.

Condevaux:2020:GGL

[16] Charles Condevaux, Stéphane Mussard, and Guillaume Zambrano. Generalized Gini linear and quadratic discriminant analyses. Metron, 78(2): 219–236, August 2020. CODEN MRONAM. ISSN 0026-1424 (print), 2281-

695X (electronic). URL https://link.springer.com/article/10.1007/
s40300-020-00178-2.

Ayad:2020:LLM

[17] Somia Ayad, Ali Laksaci, and Rachida Rouane. On the local linear modelization of the conditional density for functional and ergodic data. Metron, 78(2):237–254, August 2020. CODEN MRONAM. ISSN 0026-1424 (print),

2281-695X (electronic). URL https://link.springer.com/article/10.
1007/s40300-020-00174-6.

Fouskakis:2020:BLM

[18] D. Fouskakis, G. Petrakos, and I. Rotous. A Bayesian longitudinal model for quantifying students’ preferences regarding teaching quality indicators. Metron, 78(2):255–270, August 2020. CODEN MRONAM. ISSN 0026-1424 (print), 2281-695X (electronic). URL https://link.springer.com/

article/10.1007/s40300-020-00175-5.

Gasparini:2020:DCF

[19] Mauro Gasparini and Lidia Sacchetto. On the definition of a concentration function relevant to the ROC curve. Metron, 78(3):271–277, December 2020. CODEN MRONAM. ISSN 0026-1424 (print), 2281-695X (elec-

tronic). URL https://link.springer.com/article/10.1007/s40300-
020-00191-5.

Hoffmann:2020:SIM

[20] Rodolfo Hoffmann and Diego Camargo Botassio. Sensitivity of inequality measures considering regressive transfers with fixed relative income distance. Metron, 78(3):279–296, December 2020. CODEN MRONAM. ISSN 0026-1424 (print), 2281-695X (electronic). URL https://link.springer.

com/article/10.1007/s40300-020-00189-z.

Lando:2020:SDR

[21] Tommaso Lando and Lucio Bertoli-Barsotti. Stochastic dominance relations for generalised parametric distributions obtained through composition. Metron, 78(3):297–311, December 2020. CODEN MRONAM. ISSN 0026-1424 (print), 2281-695X (electronic). URL https://link.springer.

com/article/10.1007/s40300-020-00184-4.

Alhamzawi:2020:BRP

[22] Rahim Alhamzawi and Haithem Taha Mohammad Ali. Brq: an R package for Bayesian quantile regression. Metron, 78(3):313–328, December 2020. CODEN MRONAM. ISSN 0026-1424 (print), 2281-695X (electronic). URL

https://link.springer.com/article/10.1007/s40300-020-00190-6.

Amorim:2020:CNL

[23] Erick da Conceição Amorim and Vinícius Diniz Mayrink. Clustering nonlinear interactions in factor analysis. Metron, 78(3):329–352, December 2020. CODEN MRONAM. ISSN 0026-1424 (print), 2281-695X (elec-

tronic). URL https://link.springer.com/article/10.1007/s40300-
020-00186-2.

Rayner:2020:KWT

[24] J. C. W. Rayner and Glen Livingston Jr. The Kruskal–Wallis tests are Cochran–Mantel–Haenszel mean score tests. Metron, 78(3):353–360, December 2020. CODEN MRONAM. ISSN 0026-1424 (print), 2281-

695X (electronic). URL https://link.springer.com/article/10.1007/
s40300-020-00192-4.

Agiwal:2020:BET

[25] Varun Agiwal and Jitendra Kumar. Bayesian estimation for threshold autoregressive model with multiple structural breaks. Metron, 78(3):361– 382, December 2020. CODEN MRONAM. ISSN 0026-1424 (print), 2281-

695X (electronic). URL https://link.springer.com/article/10.1007/
s40300-020-00188-0.

Mukhopadhyay:2020:GAS

[26] Nitis Mukhopadhyay and Srawan Kumar Bishnoi. On general asymptotically second-order efficient purely sequential fixed-width confidence interval (FWCI) and minimum risk point estimation (MRPE) strategies for a normal mean and optimality. Metron, 78(3):383–409, December 2020. CODEN MRONAM. ISSN 0026-1424 (print), 2281-695X (electronic). URL https:

//link.springer.com/article/10.1007/s40300-020-00187-1.

Bowen:2021:DPD

[27] Claire McKay Bowen, Fang Liu, and Bingyue Su. Differentially private data release via statistical election to partition sequentially. Metron, 79 (1):1–31, April 2021. CODEN MRONAM. ISSN 0026-1424 (print), 2281-

695X (electronic). URL https://link.springer.com/article/10.1007/
s40300-021-00201-0.

Bourguignon:2021:NRM

[28] Marcelo Bourguignon, Manoel Santos-Neto, and Mário de Castro. A new regression model for positive random variables with skewed and long tail. Metron, 79(1):33–55, April 2021. CODEN MRONAM. ISSN 0026-1424 (print), 2281-695X (electronic). URL https://link.springer.com/

article/10.1007/s40300-021-00203-y.

Tomal:2021:ICS

[29] Jabed Tomal, Saeed Rahmati, and Ehsan Ahmed. The impact of COVID-19 on students’ marks: A Bayesian hierarchical modeling approach. Metron, 79(1):57–91, April 2021. CODEN MRONAM. ISSN 0026-1424 (print),

2281-695X (electronic). URL https://link.springer.com/article/10.
1007/s40300-021-00200-1.

Pal:2021:PAA

[30] Sanghamitra Pal and Dipika Patra. A prediction approach in adaptive sampling. Metron, 79(1):93–108, April 2021. CODEN MRONAM. ISSN 0026-1424 (print), 2281-695X (electronic). URL https://link.springer.

com/article/10.1007/s40300-020-00195-1.

Zezula:2021:DLS

[31] Ivan Zezula and Daniel Klein. On drawbacks of least squares Lehmann– Scheffé estimation of variance components. Metron, 79(1):109–119, April 2021. CODEN MRONAM. ISSN 0026-1424 (print), 2281-695X (elec-

tronic). URL https://link.springer.com/article/10.1007/s40300-
021-00196-8.

Riani:2021:ESI

[32] Marco Riani and Mia Hubert. Editorial, special issue on “Advances in Robust Statistics”. Metron, 79(2):121–125, August 2021. CODEN MRONAM. ISSN 0026-1424 (print), 2281-695X (electronic). URL https://link.

springer.com/article/10.1007/s40300-021-00213-w.

Ronchetti:2021:MCR

[33] Elvezio Ronchetti. The main contributions of robust statistics to statistical science and a new challenge. Metron, 79(2):127–135, August 2021. CODEN MRONAM. ISSN 0026-1424 (print), 2281-695X (electronic). URL https:

//link.springer.com/article/10.1007/s40300-020-00185-3.

Elsaied:2021:REN

[34] Hanan Elsaied and Roland Fried. On robust estimation of negative binomial INARCH models. Metron, 79(2):137–158, August 2021. CODEN MRONAM. ISSN 0026-1424 (print), 2281-695X (electronic). URL https:

//link.springer.com/article/10.1007/s40300-021-00207-8.

Boente:2021:RFP

[35] Graciela Boente and Matías Salibián-Barrera. Robust functional principal components for sparse longitudinal data. Metron, 79(2):159–188, August 2021. CODEN MRONAM. ISSN 0026-1424 (print), 2281-695X (elec-

tronic). URL https://link.springer.com/article/10.1007/s40300-
020-00193-3.

Ortner:2021:LPH

[36] Thomas Ortner, Peter Filzmoser, and Christian Breiteneder. Local projections for high-dimensional outlier detection. Metron, 79(2):189–206, August 2021. CODEN MRONAM. ISSN 0026-1424 (print), 2281-695X (elec-

tronic). URL https://link.springer.com/article/10.1007/s40300-
020-00183-5.

Rai:2021:REV

[37] Himanshu Rai, Sanjeev K. Tomer, and Anoop Chaturvedi. Robust estimation with variational Bayes in presence of competing risks. Metron, 79 (2):207–223, August 2021. CODEN MRONAM. ISSN 0026-1424 (print),

2281-695X (electronic). URL https://link.springer.com/article/10.
1007/s40300-021-00208-7.

Saraceno:2021:REM

[38] Giovanni Saraceno, Claudio Agostinelli, and Luca Greco. Robust estimation for multivariate wrapped models. Metron, 79(2):225–240, August 2021. CODEN MRONAM. ISSN 0026-1424 (print), 2281-695X (electronic). URL

https://link.springer.com/article/10.1007/s40300-021-00214-9.

Chakravarty:2021:IPR

[39] Satya R. Chakravarty and Palash Sarkar. An inequality paradox: relative versus absolute indices? Metron, 79(2):241–254, August 2021. CODEN MRONAM. ISSN 0026-1424 (print), 2281-695X (electronic). URL https:

//link.springer.com/article/10.1007/s40300-021-00206-9.

Giorgi:2021:SNI

[40] Giovanni Maria Giorgi. Some notes on Iddo Eliazar’s monograph “Power Laws: A Statistical trek”. Metron, 79(2):255–257, August 2021. CODEN MRONAM. ISSN 0026-1424 (print), 2281-695X (electronic). URL https:

//link.springer.com/article/10.1007/s40300-021-00205-w.

Bohning:2021:IHM

[41] Dankmar Böhning, Heinz Holling, and Patarawan Sangnawakij. Investigating heterogeneity in meta-analysis of studies with rare events. Metron, 79 (3):259–272, December 2021. CODEN MRONAM. ISSN 0026-1424 (print),

2281-695X (electronic). URL https://link.springer.com/article/10.
1007/s40300-021-00211-y.

Ando:2021:DMR

[42] Shuji Ando. A directional measure for reverse global symmetry in square contingency tables with application to unaided vision data. Metron, 79(3): 273–283, December 2021. CODEN MRONAM. ISSN 0026-1424 (print),

2281-695X (electronic). URL https://link.springer.com/article/10.
1007/s40300-021-00220-x.

Fajardo:2021:BEF

[43] Jesús Fajardo and Pedro Harmath. Boundary estimation with the fuzzy set density estimator. Metron, 79(3):285–302, December 2021. CODEN MRONAM. ISSN 0026-1424 (print), 2281-695X (electronic). URL https:

//link.springer.com/article/10.1007/s40300-021-00210-z.

Bamba:2021:MES

[44] Lanciné Bamba, Ouagnina Hili, and Assi N’Guessan. M-estimate for the stationary hyperbolic GARCH models. Metron, 79(3):303–351, December 2021. CODEN MRONAM. ISSN 0026-1424 (print), 2281-695X (elec-

tronic). URL https://link.springer.com/article/10.1007/s40300-
021-00221-w.

Liu:2021:MET

[45] Yan Liu and Masanobu Taniguchi. Minimax estimation for time series models. Metron, 79(3):353–359, December 2021. CODEN MRONAM. ISSN 0026-1424 (print), 2281-695X (electronic). URL https://link.springer.

com/article/10.1007/s40300-021-00217-6.

Tomal:2021:BPL

[46] Jabed H. Tomal and Hafizur Rahman. A Bayesian piecewise linear model for the detection of breakpoints in housing prices. Metron, 79(3):361– 381, December 2021. CODEN MRONAM. ISSN 0026-1424 (print), 2281-

695X (electronic). URL https://link.springer.com/article/10.1007/
s40300-021-00223-8.

Azimi:2021:LEC

[47] Reza Azimi and Mahdy Esmailian. Letter to the editor: Correction to “Lindley-exponential distribution: properties and applications”. Metron, 79(3):383–386, December 2021. CODEN MRONAM. ISSN 0026-1424 (print), 2281-695X (electronic). URL https://link.springer.com/

article/10.1007/s40300-021-00212-x.

Mosler:2021:CCU

[48] Karl Mosler. Correction to: Commentary on “From unidimensional to multidimensional inequality: a review”. Metron, 79(3):387, December 2021. CODEN MRONAM. ISSN 0026-1424 (print), 2281-695X (elec-

tronic). URL https://link.springer.com/article/10.1007/s40300-
021-00219-4. See [8].

Anonymous:2021:CNM

[49] Anonymous. Considerations on Nitis Mukhopadhyay and Partha Pratim Sengupta’s book “Gini inequality index: methods and applications”.

Metron, 79(3):389–391, December 2021. CODEN MRONAM. ISSN 0026-1424 (print), 2281-695X (electronic). URL https://link.springer.com/

article/10.1007/s40300-021-00222-9.

Alfo:2022:GMG

[50] Marco Alfo, Alessio Guandalini, and Roberto Zelli. Giovanni Maria Giorgi (1947–2021) obituary. Metron, 80(1):1–2, April 2022. CODEN MRONAM. ISSN 0026-1424 (print), 2281-695X (electronic). URL https://link.

springer.com/article/10.1007/s40300-022-00232-1.

Bini:2022:SES

[51] Matilde Bini, Marcello Chiodi, and Paola Zuccolotto. Statistical evaluation systems at 360^{∘}: techniques, technologies and new frontiers. Metron, 80 (1):3–8, April 2022. CODEN MRONAM. ISSN 0026-1424 (print), 2281-

695X (electronic). URL https://link.springer.com/article/10.1007/
s40300-022-00233-0.

Camminatiello:2022:ATW

[52] Ida Camminatiello, Antonello D’Ambra, and Luigi D’Ambra. The association in two-way ordinal contingency tables through global odds ratios. Metron, 80(1):9–22, April 2022. CODEN MRONAM. ISSN 0026-1424 (print), 2281-695X (electronic). URL https://link.springer.com/

article/10.1007/s40300-021-00224-7.

Cavicchia:2022:MGH

[53] Carlo Cavicchia and Pasquale Sarnacchiaro. A multi-group higher-order factor analysis for studying the gender-effect in Teacher Job Satisfaction. Metron, 80(1):23–38, April 2022. CODEN MRONAM. ISSN 0026-1424 (print), 2281-695X (electronic). URL https://link.springer.com/

article/10.1007/s40300-021-00204-x. See correction [54].

Cavicchia:2022:CMG

[54] Carlo Cavicchia and Pasquale Sarnacchiaro. Correction to: A multi-group higher-order factor analysis for studying the gender-effect in Teacher Job Satisfaction. Metron, 80(1):39, April 2022. CODEN MRONAM. ISSN 0026-1424 (print), 2281-695X (electronic). URL https://link.springer.

com/article/10.1007/s40300-021-00209-6. See [53].

DAmbra:2022:GTS

[55] Antonello D’Ambra, Pietro Amenta, and Antonio Lucadamo. The generalized Taguchi’s statistic: a passenger satisfaction evaluation. Metron, 80(1):41–60, April 2022. CODEN MRONAM. ISSN 0026-1424 (print),

2281-695X (electronic). URL https://link.springer.com/article/10.
1007/s40300-021-00202-z.

Iannario:2022:MCP

[56] Maria Iannario and Anna Clara Monti. Modelling consumer perceptions of service quality for urban public transport systems using statistical models for ordinal data. Metron, 80(1):61–76, April 2022. CODEN MRONAM. ISSN 0026-1424 (print), 2281-695X (electronic). URL https://link.

springer.com/article/10.1007/s40300-021-00197-7.

Iaco:2022:UMM

[57] Sandra De Iaco and Sabrina Maggio. Using multilevel models to evaluate the attitude of separate waste collection in young people. Metron, 80(1): 77–95, April 2022. CODEN MRONAM. ISSN 0026-1424 (print), 2281-

695X (electronic). URL https://link.springer.com/article/10.1007/
s40300-020-00194-2.

Pagani:2022:EPP

[58] Laura Pagani and Demetrio Panarello. Evaluation of a program for promoting physical activity and well-being: Friuli Venezia Giulia in Movimento. Metron, 80(1):97–120, April 2022. CODEN MRONAM. ISSN 0026-1424 (print), 2281-695X (electronic). URL https://link.springer.

com/article/10.1007/s40300-021-00215-8.

Biffignandi:2022:BPA

[59] Silvia Biffignandi and Alessandro Zeli. Building panels from archives: the longitudinal representativity. Metron, 80(1):121–138, April 2022. CODEN MRONAM. ISSN 0026-1424 (print), 2281-695X (electronic). URL https:

//link.springer.com/article/10.1007/s40300-021-00199-5.

Mariani:2022:MVB

[60] Paolo Mariani and Andrea Marletta. Missing value or behaviour: how to increase the signal of social media data. Metron, 80(2):139–151, August 2022. CODEN MRONAM. ISSN 0026-1424 (print), 2281-695X (electronic). URL

https://link.springer.com/article/10.1007/s40300-021-00216-7.

Davino:2022:HMQ

[61] C. Davino, R. Romano, and D. Vistocco. Handling multicollinearity in quantile regression through the use of principal component regression. Metron, 80(2):153–174, August 2022. CODEN MRONAM. ISSN 0026-1424 (print), 2281-695X (electronic). URL https://link.springer.com/

article/10.1007/s40300-022-00230-3.

Vanacore:2022:WDM

[62] Amalia Vanacore and Maria Sole Pellegrino. A weighted distance metric for assessing ranking dissimilarity and inter-group heterogeneity. Metron, 80(2):175–185, August 2022. CODEN MRONAM. ISSN 0026-1424 (print),

2281-695X (electronic). URL https://link.springer.com/article/10.
1007/s40300-021-00198-6.

Giacalone:2022:OFA

[63] Massimiliano Giacalone. Optimal forecasting accuracy using Lp-norm combination. Metron, 80(2):187–230, August 2022. CODEN MRONAM. ISSN 0026-1424 (print), 2281-695X (electronic). URL https://link.springer.

com/article/10.1007/s40300-021-00218-5.

Galarza:2022:EAE

[64] Christian E. Galarza, Larissa A. Matos, and Victor H. Lachos. An EM algorithm for estimating the parameters of the multivariate skew-normal distribution with censored responses. Metron, 80(2):231–253, August 2022. CODEN MRONAM. ISSN 0026-1424 (print), 2281-695X (electronic). URL

https://link.springer.com/article/10.1007/s40300-021-00227-4.

Sefidi:2022:AOC

[65] Saeide Sefidi, Mojtaba Ganjali, and Taban Baghfalaki. Analysis of ordinal and continuous longitudinal responses using pair copula construction. Metron, 80(2):255–280, August 2022. CODEN MRONAM. ISSN 0026-1424 (print), 2281-695X (electronic). URL https://link.springer.com/

article/10.1007/s40300-022-00231-2.

Ahanda:2022:RLC

[66] Benoit Ahanda, Daniel E. Osborne, and Leif Ellingson. Robustness of lognormal confidence regions for means of symmetric positive definite matrices when applied to mixtures of lognormal distributions. Metron, 80(3): 281–303, December 2022. CODEN MRONAM. ISSN 0026-1424 (print),

2281-695X (electronic). URL https://link.springer.com/article/10.
1007/s40300-022-00234-z.

Turetsky:2022:ORV

[67] Vladimir Turetsky and Emil Bashkansky. Ordinal response variation of the polytomous Rasch model. Metron, 80(3):305–330, December 2022. CODEN MRONAM. ISSN 0026-1424 (print), 2281-695X (electronic). URL https:

//link.springer.com/article/10.1007/s40300-022-00229-w.

Sisuma:2022:NPT

[68] M. S. Sisuma and P. G. Sankaran. Non-parametric test of recurrent cumulative incidence functions for competing risks models. Metron, 80(3): 331–342, December 2022. CODEN MRONAM. ISSN 0026-1424 (print),

2281-695X (electronic). URL https://link.springer.com/article/10.
1007/s40300-022-00228-x.

Lau:2022:PPD

[69] F. Din-Houn Lau and Sebastian Krumscheid. Plateau proposal distributions for adaptive component-wise multiple-try Metropolis. Metron, 80(3): 343–370, December 2022. CODEN MRONAM. ISSN 0026-1424 (print),

2281-695X (electronic). URL https://link.springer.com/article/10.
1007/s40300-022-00235-y.

Castro:2022:FCC

[70] Tuany de Paula Castro, Carlos Daniel Paulino, and Julio M. Singer. A fair comparison of credible and confidence intervals: an example with binomial proportions. Metron, 80(3):371–382, December 2022. CODEN MRONAM. ISSN 0026-1424 (print), 2281-695X (electronic). URL https:

//link.springer.com/article/10.1007/s40300-021-00225-6.

Ahn:2022:CCR

[71] Soohyun Ahn, Xinlei Wang, Mumu Wang, and Johan Lim. On continuity correction for RSS-structured cluster randomized designs with binary outcomes. Metron, 80(3):383–397, December 2022. CODEN MRONAM. ISSN 0026-1424 (print), 2281-695X (electronic). URL https://link.springer.

com/article/10.1007/s40300-021-00226-5.

Ranalli:2023:FSI

[72] M. Giovanna Ranalli, Jean-François Beaumont, Gaia Bertarelli, and Nathalie Shlomo. Foreword to the special issue on “Survey Methods for Statistical Data Integration and New Data Sources”. Metron, 81(1):1–3, April 2023. CODEN MRONAM. ISSN 0026-1424 (print), 2281-695X (elec-

tronic). URL https://link.springer.com/article/10.1007/s40300-
023-00248-1.

Quatember:2023:EIC

[73] Andreas Quatember. Efficient item count techniques with one or two lists. Metron, 81(1):5–19, April 2023. CODEN MRONAM. ISSN 0026-1424 (print), 2281-695X (electronic). URL https://link.springer.com/

article/10.1007/s40300-023-00240-9.

Fellows:2023:MNP

[74] Ian E. Fellows and Mark S. Handcock. Modeling of networked populations when data is sampled or missing. Metron, 81(1):21–35, April 2023. CODEN MRONAM. ISSN 0026-1424 (print), 2281-695X (electronic). URL https:

//link.springer.com/article/10.1007/s40300-023-00246-3.

Varriale:2023:MSS

[75] Roberta Varriale and Marco Alfo’. Multi-source statistics on employment status in Italy, a machine learning approach. Metron, 81(1):37–63, April 2023. CODEN MRONAM. ISSN 0026-1424 (print), 2281-695X (elec-

tronic). URL https://link.springer.com/article/10.1007/s40300-
023-00242-7.

Huang:2023:DFA

[76] Chien-Min Huang and F. Jay Breidt. A dual-frame approach for estimation with respondent-driven samples. Metron, 81(1):65–81, April 2023. CODEN MRONAM. ISSN 0026-1424 (print), 2281-695X (electronic). URL https:

//link.springer.com/article/10.1007/s40300-023-00241-8.

Salvatore:2023:INP

[77] Camilla Salvatore. Inference with non-probability samples and survey data integration: a science mapping study. Metron, 81(1):83–107, April 2023. CODEN MRONAM. ISSN 0026-1424 (print), 2281-695X (electronic). URL

https://link.springer.com/article/10.1007/s40300-023-00243-6.

Solari:2023:SFF

[78] Fabrizio Solari, Antonella Bernardini, and Nicoletta Cibella. Statistical framework for fully register based population counts. Metron, 81(1):109– 129, April 2023. CODEN MRONAM. ISSN 0026-1424 (print), 2281-

695X (electronic). URL https://link.springer.com/article/10.1007/
s40300-023-00244-5.

Severini:2023:ILI

[79] Thomas A. Severini. Integrated likelihood inference in multinomial distributions. Metron, 81(2):131–142, August 2023. CODEN MRONAM. ISSN 0026-1424 (print), 2281-695X (electronic). URL https://link.springer.

com/article/10.1007/s40300-022-00236-x.

Persson:2023:TEP

[80] Rasmus A. X. Persson. Theoretical evaluation of partial credit scoring of the multiple-choice test item. Metron, 81(2):143–161, August 2023. CODEN MRONAM. ISSN 0026-1424 (print), 2281-695X (electronic). URL https:

//link.springer.com/article/10.1007/s40300-022-00237-w.

Trick:2023:PEB

[81] Susanne Trick, Constantin A. Rothkopf, and Frank Jäkel. Parameter estimation for a bivariate beta distribution with arbitrary beta marginals and positive correlation. Metron, 81(2):163–180, August 2023. CODEN MRONAM. ISSN 0026-1424 (print), 2281-695X (electronic). URL https:

//link.springer.com/article/10.1007/s40300-023-00247-2.

Chen:2023:MGP

[82] Zexun Chen, Jun Fan, and Kuo Wang. Multivariate Gaussian processes: definitions, examples and applications. Metron, 81(2):181–191, August 2023. CODEN MRONAM. ISSN 0026-1424 (print), 2281-695X (elec-

tronic). URL https://link.springer.com/article/10.1007/s40300-
023-00238-3.

Mai:2023:SCB

[83] The Tien Mai. Simulation comparisons between Bayesian and de-biased estimators in low-rank matrix completion. Metron, 81(2):193–214, August 2023. CODEN MRONAM. ISSN 0026-1424 (print), 2281-695X (elec-

tronic). URL https://link.springer.com/article/10.1007/s40300-
023-00239-2.

Ndwandwe:2023:TPT

[84] L. Ndwandwe, J. S. Allison, L. Santana, and I. J. H. Visagie. Testing for the Pareto type I distribution: a comparative study. Metron, 81(2): 215–256, August 2023. CODEN MRONAM. ISSN 0026-1424 (print), 2281-

695X (electronic). URL https://link.springer.com/article/10.1007/
s40300-023-00252-5.

Perri:2023:ESI

[85] Pier Francesco Perri, Gennaro Punzo, and Carlo Cavicchia. Editorial to the special issue on “Advances in statistical modelling, methods and applications”. Metron, 81(3):257–258, December 2023. CODEN MRONAM. ISSN 0026-1424 (print), 2281-695X (electronic). URL https://link.springer.

com/article/10.1007/s40300-023-00260-5.

Maricic:2023:IUC

[86] Milica Maricic and Veljko Jeremic. Imposing unsupervised constraints to the Benefit-of-the-Doubt (BoD) model. Metron, 81(3):259–296, December 2023. CODEN MRONAM. ISSN 0026-1424 (print), 2281-695X (elec-

tronic). URL https://link.springer.com/article/10.1007/s40300-
023-00254-3.

Ren:2023:GIW

[87] Yijun Ren, Alessandra Guglielmi, and Lara Maestripieri. Gender inequalities at work in southern Europe. Metron, 81(3):297–322, December 2023. CODEN MRONAM. ISSN 0026-1424 (print), 2281-695X (electronic). URL

https://link.springer.com/article/10.1007/s40300-023-00245-4.

Simone:2023:DMP

[88] Rosaria Simone, Marcella Corduas, and Domenico Piccolo. Dynamic modelling of price expectations and judgments. Metron, 81(3):323–342, December 2023. CODEN MRONAM. ISSN 0026-1424 (print), 2281-695X (elec-

tronic). URL https://link.springer.com/article/10.1007/s40300-
023-00251-6.

Musella:2023:ESH

[89] Gaetano Musella, Rosalia Castellano, and Emma Bruno. Evaluating the spatial heterogeneity of innovation drivers: a comparison between GWR and GWPR. Metron, 81(3):343–365, December 2023. CODEN MRONAM. ISSN 0026-1424 (print), 2281-695X (electronic). URL https://link.

springer.com/article/10.1007/s40300-023-00249-0.

Alboni:2023:STR

[90] Fabrizio Alboni, Pasquale Pavone, and Margherita Russo. The search for topics related to electric mobility: a comparative analysis of some of the most widely used methods in the literature. Metron, 81(3):367–391, December 2023. CODEN MRONAM. ISSN 0026-1424 (print), 2281-695X (elec-

tronic). URL https://link.springer.com/article/10.1007/s40300-
023-00255-2.

Ranalli:2024:FSI

[91] M. Giovanna Ranalli, Jean-François Beaumont, Gaia Bertarelli, and Natalie Shlomo. Foreword to the special issue on “Survey Methods for Statistical Data Integration and New Data Sources: tools and real data applications for official statistics”. Metron, 82(1):1–3, April 2024. CODEN MRONAM. ISSN 0026-1424 (print), 2281-695X (electronic). URL https:

//link.springer.com/article/10.1007/s40300-024-00270-x. See cor-

rection [98].

Bernardini:2024:IDI

[92] Antonella Bernardini, Angela Chieppa, Nicoletta Cibella, and Donatella Zindato. Improving the design of the Italian permanent population and housing census: a transition towards a massive use of administrative data. Metron, 82(1):5–17, April 2024. CODEN MRONAM. ISSN 0026-1424 (print), 2281-695X (electronic). URL https://link.springer.com/

article/10.1007/s40300-023-00256-1.

Vitiis:2024:ASD

[93] Claudia De Vitiis, Stefano Falorsi, Alessio Guandalini, Francesca Inglese, Paolo Righi, and Marco D. Terribili. Adaptive sampling design for the Italian social sample surveys: an application on the population census. Metron, 82(1):19–35, April 2024. CODEN MRONAM. ISSN 0026-1424 (print),

2281-695X (electronic). URL https://link.springer.com/article/10.
1007/s40300-023-00262-3.

Corazziari:2024:ITG

[94] Isabella Corazziari, Gabriele Ascari, and Maria Giuseppina Muratore. If the tools to gather information affect data quality: violence against women survey case. Metron, 82(1):37–70, April 2024. CODEN MRONAM. ISSN 0026-1424 (print), 2281-695X (electronic). URL https://link.springer.

com/article/10.1007/s40300-024-00266-7.

Salvatore:2024:ABS

[95] Camilla Salvatore, Silvia Biffignandi, and Annamaria Bianchi. Augmenting business statistics information by combining traditional data with textual data: a composite indicator approach. Metron, 82(1):71–91, April 2024. CODEN MRONAM. ISSN 0026-1424 (print), 2281-695X (electronic). URL

https://link.springer.com/article/10.1007/s40300-023-00261-4.

DAlo:2024:EMM

[96] Michele D’Aló, Stefano Falorsi, and Andrea Fasulo. Mind, a methodology for multivariate small area estimation with multiple random effects. Metron, 82(1):93–107, April 2024. CODEN MRONAM. ISSN 0026-1424 (print), 2281-695X (electronic). URL https://link.springer.com/

article/10.1007/s40300-023-00258-z. See publisher correction [99].

Marcis:2024:HSV

[97] Laura Marcis, Maria Chiara Pagliarella, and Renato Salvatore. How the sampling variances affect the linear predictor of the Fay–Herriot model. Metron, 82(1):109–130, April 2024. CODEN MRONAM. ISSN 0026-1424 (print), 2281-695X (electronic). URL https://link.springer.com/

article/10.1007/s40300-023-00250-7.

Ranalli:2024:CFS

[98] M. Giovanna Ranalli, Jean-François Beaumont, Gaia Bertarelli, and Natalie Shlomo. Correction to: Foreword to the special issue on “Survey Methods for Statistical Data Integration and New Data Sources”. Metron, 82(1):131, April 2024. CODEN MRONAM. ISSN 0026-1424 (print), 2281-

695X (electronic). URL https://link.springer.com/article/10.1007/
s40300-024-00268-5. See [91].

DAlo:2024:PCM

[99] Michele D’Aló, Stefano Falorsi, and Andrea Fasulo. Publisher correction: Mind, a methodology for multivariate small area estimation with multiple random effects. Metron, 82(1):133, April 2024. CODEN MRONAM. ISSN 0026-1424 (print), 2281-695X (electronic). URL https://link.springer.

com/article/10.1007/s40300-024-00265-8. See [96].